OpenStudy (anonymous):

An equation is shown below: 2 over 3 multiplied by x plus 8 over 3 equals 2 to the power of x What is the solution to the equation? x = –2 x = –1 x = 1 x = 2

sammixboo (sammixboo):

Does your problem look like this? \(\Large \frac{2}{3} \times x~+~\frac{8}{3}=~2^x\)

OpenStudy (anonymous):

thanks you so much when i need help can i tag u